OpenStudy (anonymous):

What is a solution of x2 - 7x = -5? (Quadratics)

OpenStudy (bohotness):

x =(7-√29)/2= 0.807 x =(7+√29)/2= 6.193

OpenStudy (studygurl14):

@jon198 do you know how to use the quadratic formula to solve quadratic equations?

OpenStudy (anonymous):

somewhat I know the formula but I keep getting weird answers @StudyGurl14

OpenStudy (studygurl14):

Okay. hold on a sec while i type it up, ok?

OpenStudy (studygurl14):

\(\large x^2-7x=-5\rightarrow x^2-7x+5=0\) With a quadratic equation in the form \(\large ax^2+bx+c=0\), \(\large x=\Large\frac{-b\pm\sqrt{b^2-4ac}}{2a}\)

OpenStudy (studygurl14):

What do a, b, and c equal?

OpenStudy (anonymous):

wait is it 7 +/- Sqrt of 29/ 2?

OpenStudy (studygurl14):

Yes, actually. :)

OpenStudy (anonymous):

Boo yah thanks!